OpenStudy (stormoshawty):

A hotel offers two activity packages. One costs $192 and includes 3 hrs of horseback riding and 2 hrs of para-sailing. The second costs $213 and includes 2 hrs of horseback riding and 3 hrs of para-sailing. What is the cost for 1 hr of each activity?

OpenStudy (stormoshawty):

@OwlCoffee

OpenStudy (owlcoffee):

okay, so this is clearly a problem where we have to use a system of equations. so let's say the hours of horseback riding is "m" and the para sailing hours is "n". The first package says that it costs $192 and includes 3 hours of "m" and 2 hours of "n". The second one costs $213 but includes 2 hours of "m" and 3 hours of "n". Let's tradcue them into two equations: \[3m+2n=192\] \[2m+3n=213\] Sine we've been using elimination to do the other two, let's use that for this one. I'll multiply the first equation by -3 and the second by 2: \[-9m-6n=-576\] \[4m+6n=426\] Now, let's sum the equations, reorder and operate the variables with constants: \[-5m=-150\] So, solving for "m" we get that m=30. Let's replace it on the initial equations: \[3(30)+2n=192\] Sustracting 90 on each sides and operating: \[2n=102\] so we get that n=51 we can conclude that: \[m=30\] \[n=51\] Traducing it to the problem, the cost of horseback riding hours are $30 each and the cost of para-sailing hours are $51 each.